# SynthVLM: High-Efficiency and High-Quality Synthetic Data for Vision Language Models

### v1

@@ 1. Introduction
  • New Perspective. To the best of our knowledge, we are the first to utilize generated images for training VLLMs. This sets a precedent for utilizing generated images to train large-scale VLLMs, offering a solution to the limitation of data availability, quality, and privacy.
@@ 3. Method
In this section, we introduce our data generation pipeline and then compare the generated dataset with other commonly used datasets. In subsection 3.1, we explore the construction of our caption set and the subsequent generation of 1 million caption-image pairs using diffusion models. In subsection 3.2, we outline the filtering process applied to the generated 1 million image-text pairs. By utilizing CLIPScore as the alignment criterion, we meticulously select 100k high-quality, well-aligned caption-image pairs, ensuring the robustness and relevance of our dataset. In subsection 3.3, we demonstrate the effectiveness of our method by comparing the SynthVLM dataset with other existing datasets. In subsection 3.4, we summarize the VLLMs training pipeline by utilizing the synthetic dataset.
@@ 3.1. Synthetic Dataset Construction
Caption Curation. To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. This filtering process was performed using GPT-3, ensuring that only high-quality, informative captions were used for training. For the remaining captions, we calculated the CLIPScore for these captions and their corresponding raw images, selecting the top 40% of image-caption pairs with the highest scores. These selected captions were included in the candidate caption set. At this step, we believe that captions with better image-caption alignment are more likely to generate high-quality, well-aligned images. Ultimately, we sampled a dataset of 1,000k captions for data generation. By using only captions, our method significantly reduces storage overhead and processing time. The caption curation pipeline is summarized in Figure 3(a).
3.2. Synthetic Data Selection
In this section, we introduce the quality control process for synthetic datasets. To obtain high-quality image-text pairs, we leverage CLIPScore, a metric that measures the cosine similarity between images and their corresponding captions, effectively assessing the alignment between image and text.
@@ 3.2. Synthetic Data Selection
As shown in Figure 3(b), we initially computed CLIPScores for the 1, [^0] K synthetic image-caption pairs. We then selected the top 100K pairs that demonstrated the highest scores, indicating the most accurate and meaningful matches between images and captions. By curating this subset, we constructed a high-quality, highly aligned synthetic dataset, ensuring we retain the pairs with the highest CLIPScores.
@@ 4.2. Synthetic Data Achieves SoTA Performance
These results demonstrate the strong alignment capability of our synthetic data and provide a new paradigm for an effective visual understanding of model modality alignment using generated data. During pre-training, it is common to train on all available data due to uncertainty about data selection. Here, we offer 100k high-quality synthetic data as a benchmark for selecting aligned generated data efficiently.
@@ 4.3. Effective Vision Language Alignment
From Table 5, it is evident that our Synth Select 100k model outperforms the Baseline on all MMLU benchmarks. Additionally, as shown in Table 4, our model surpasses the baseline on SQA benchmarks.
@@
Table 6. Comparison of data point utilization and computational costs for data transfer. This indicates that our SynthVLM have superior efficiency compared to other methods
Models SynthVLM LLaVA w/o selction
@@ 4.4. Efficient Vision Language Alignment
To address Q3, we first compare the data usage and computational resource usage between the baseline, unselected data, and the synthetic dataset. Next, we compare the transfer usage of the synthetic 100k data and the llava 558k data.
Data Usage and Computational Resource Usage
As shown in Figure 5, with the data quality selection module, we use only 19% of the llava data and 10% of the synthetic data while achieving state-of-the-art performance. This indicates that our data selection method can reduce computational usage by more than 80%.
Data Transfer Efficiency
Besides the efficient use of data and computational resources, utilizing only captions to generate images can significantly reduce data transfer usage. For instance, transferring 1,000,000 captions requires only 330 MB, while transferring 100,000 generated captions needs just 33 MB. In contrast, transferring 1,000,000 image captions would normally exceed 50 GB in data storage. From Table 6 we can see that transfer LLaVA pre-train cost 27G, which is very costly. This demonstrates that using captions alone is far more efficient than transmitting image captions.
Overall, from the perspectives of data usage, computational resource usage, and data transfer efficiency, our synthetic data can efficiently align image and text modalities, indicating a strong potential for efficient modality alignment.
@@ 4.5. Privacy Protection Pre-training
Table 7. An ablation study of visual understanding ability was conducted. We first removed the data generation model and then the data selection model. The performance decreased when either module was removed.
@@
w/o selection 100k Vicuna-1.5-7B 69.9$\downarrow$ 67.7$\downarrow$ 30.2$\downarrow$ 50.2$\uparrow$ 79.1$\downarrow$ 62.2$\downarrow$ 63.5$\downarrow$ 1421.7$\downarrow$ 301.8$\downarrow$ 87.3$\uparrow$ 40.6$\downarrow$
@@
w/o selection 100k Vicuna-1.5-13B 74.1$\downarrow$ 70.5$\downarrow$ 35.6$\uparrow$ 53.2$\downarrow$ 79.7$\downarrow$ 63.1$\downarrow$ 67.5$\downarrow$ 1512.7$\downarrow$ 303.2$\downarrow$ 86.9$\downarrow$ 53.0$\downarrow$
Table 8. An ablation study of modality alignment was conducted. We first removed the data generation model, followed by the data selection model. The performance decreased when either module was removed.
@@ 4.6. Ablation Study
When excluding the data generation module, the performance of the model decreases across all benchmarks. As shown in Table 7 and Table 8, ”w/o generation 100k” represents the model without the data generation module. This version exhibits lower accuracy in all benchmarks compared to the full SynthVLM model. This highlights the importance of data generation in maintaining high performance.
@@
Similarly, when the data selection module is excluded, as represented by ”w/o selection 100k” in Table 7 and Table 8. The model’s performance also declines in most benchmarks. Exclusing data selection module version exhibits lower accuracy in all benchmarks compared to the full SynthVLM model. This underscores the critical role of the data selection process in achieving optimal results.

### v2

@@ 1. Introduction
  • New Perspective. To the best of our knowledge, we are the first to utilize Caption-to-Image strategy to construct a high-aligned training dataset for VLLMs and achieved superior performance with just 100K data points(only 18% of the official dataset size). This sets a precedent for utilizing generated images to train large-scale VLLMs, offering a solution to the limitation of data availability, quality, and privacy.
@@ 3. Method
In this section, we introduce our data generation pipeline and then compare the generated dataset with other commonly used datasets. In subsection 3.1, we explore the construction of our caption set and the subsequent generation of 1 million caption-image pairs using diffusion models. In subsection 3.2, we outline the filtering process applied to the generated 1 million image-text pairs. Then we meticulously select 100k high-quality, well-aligned caption-image pairs, ensuring the robustness and relevance of our dataset. In subsection 3.3, we demonstrate the effectiveness of our method by comparing the SynthVLM dataset with other existing datasets. In subsection 3.4, we summarize the VLLMs training pipeline by utilizing the synthetic dataset.
@@ 3.1. Synthetic Dataset Construction
Caption Curation. To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. This filtering process was performed using GPT-3, ensuring that only high-quality, informative captions were used for training. For the remaining captions, we calculated the CLIPScore for these captions and their corresponding raw images. CLIPScore is a metric that measures the cosine similarity between images and their corresponding captions.
@@
We selected the top 40% of image-caption pairs with the highest scores. These selected captions were included in the candidate caption set. At this step, we believe that captions with better image-caption alignment are more likely to generate high-quality, well-aligned images. Ultimately, we sampled a dataset of 1,000k captions for data generation. By using only captions, our method significantly reduces storage overhead and processing time. The caption curation pipeline is summarized in Figure 3(a).
3.2. Synthetic Data Selection
In this section, we introduce the quality control process for synthetic datasets. To further ensure the alignment between images and their corresponding text descriptions, we employ CLIPScore a second time, evaluating the quality of image-text pairs with enhanced precision.
As shown in Figure 3(b), we initially computed CLIPScores for the 1, [^0] K synthetic image-caption pairs. We then selected the top 100K pairs that demonstrated the highest scores, indicating the most accurate and meaningful matches between images and captions. By curating this subset, we constructed a high-quality, highly aligned synthetic dataset.
@@ 4.3. Effective Vision Language Alignment
From Table 5, it is evident that our Synth Select 100k model outperforms the Baseline on all MMLU benchmarks and SQA benchmarks.
@@
Table 6. Comparison of data utilization for generating image-caption pairs. This indicates that our SynthVLM have superior efficiency compared to other methods.
Methods SynthVLM LLaVA w/o selection
@@ 4.4. Efficient Vision Language Alignment
To address Q3, we examine the computational resource usage during training and evaluate data utilization efficiency for generating image-caption pairs.
Computational Resource Usage
As shown in Table 6, by integrating a data selection module, our approach utilizes only 19% of the LLAVA data and 10% of the original synthetic data while achieving SoTA performance. This demonstrates that our data selection method can reduce computational usage by more than 80%.
Data Utilization Efficiency
Table 6 compares the data utilization of our proposed method with the conventional BLIP-LCS method, which employs the BLIP model for caption generation. Our approach is significantly more efficient, requiring only 330 MB for captions to generate 1,000,000 image-caption pairs, compared to the traditional methods, which may exceed 50 GB for images. This substantial difference highlights the efficiency of relying solely on captions rather than images in generating image-caption pairs.
Overall, our method efficiently aligns image and text modalities, demonstrating strong potential for effective modality alignment. Additionally, acquiring captions is considerably less resource-intensive than obtaining images. By integrating existing large language models (LLMs), it is feasible to fully automate the generation of both captions and images, further enhancing the efficiency of dataset construction in VLLMs.
@@ 4.5. Privacy Protection Pre-training
Table 7. Ablation study of visual understanding ability and pure language ability. The results demonstrate that removing either the data generation or data selection module results in a performance drop.
@@
w/o selection 100k Vicuna-1.5-7B 69.9$\downarrow$ 67.7$\downarrow$ 30.2$\downarrow$ 50.2 79.1$\downarrow$ 62.2$\downarrow$ 63.5$\downarrow$ 1421.7$\downarrow$ 301.8$\downarrow$ 87.3 40.6$\downarrow$
@@
w/o selection 100k Vicuna-1.5-13B 74.1$\downarrow$ 70.5$\downarrow$ 35.6 53.2$\downarrow$ 79.7$\downarrow$ 63.1$\downarrow$ 67.5$\downarrow$ 1512.7$\downarrow$ 303.2$\downarrow$ 86.9$\downarrow$ 53.0$\downarrow$
Table 8. Ablation study of modality alignment. The results demonstrate that removing either the data generation or data selection module results in a performance drop.
@@ 4.6. Ablation Study
The exclusion of the data generation module significantly impacts the model’s performance, as illustrated in Tables 7 and 8, labeled as ”w/o generation 100k”. The variant without this module demonstrates markedly lower accuracy across all benchmarks. These results emphasize the crucial role of the data generation process in sustaining the high performance of the SynthVLM model. This also underscores SynthVLM’s potential in constructing highly aligned datasets.
@@
The absence of the data selection module similarly leads to a noticeable decline in performance, indicated as ”w/o selection 100k” in Tables 7 and 8. Given the inherent randomness of diffusion models, which inevitably generate some low-quality images, the data selection module is crucial for removing these subpar elements.